OpenStudy (anonymous):

How many pieces of ribbon, each 1/6m long, can be cut from a ribbon 7/8m long? ___ pieces of ribbon can be cut. That is 5 whole pieces of ribbon ___ of a piece, or ___ m, left over.

OpenStudy (anonymous):

5.25

OpenStudy (anonymous):

How many m is left over?

OpenStudy (anonymous):

0.27m

OpenStudy (anonymous):

How did you find the answer of 0.27 m? Can someone explain to me so I understand? :/

OpenStudy (anonymous):

5 whole pieces, 0.25 of a piece, 0.21875m

OpenStudy (anonymous):

1/6 = 0.166 7/8 = 0.875 0.875/0.166 = 5.27

OpenStudy (anonymous):

0.25 part of 7/8 m piece . ie (7/8)*0.25=0.21875m

OpenStudy (anonymous):

Both of you have different answers and I can't really decide which one's correct.

OpenStudy (anonymous):

ksram is pretty good. both are close enough

OpenStudy (anonymous):

5 pieces of ribbon can be cut. That is 5 whole pieces of ribbon 7/8 of a piece, or 0.27 m, left over.